OpenStudy (anonymous):

sqrt5 -6sqrt45 + 2sqrt125

OpenStudy (callisto):

\[\sqrt5 -6\sqrt{45} + 2\sqrt{125}\]\[=\sqrt5 -6\sqrt{9\times 5} + 2\sqrt{25 \times 5}\]\[=\sqrt5 -6\sqrt{3^2\times 5} + 2\sqrt{5^2 \times 5}\]\[=\sqrt5 -6(3)\sqrt{5} + 2(5)\sqrt{5}\]\[=\sqrt5 -18\sqrt{5} + 10\sqrt{5}\]\[=...?\]

OpenStudy (anonymous):

This is what I have but didn't know what to do about the sqrt5

OpenStudy (anonymous):

-7sqrt5 Thanks.

OpenStudy (callisto):

Welcome...